OpenStudy (anonymous):

The ratio of the ages of Mandy and Sandy is 2:5. After 8 years, their ages will be in the ratio 1:2. What is the difference between their present ages? I think the answer is 24 but I am not sure.

OpenStudy (anonymous):

24 is correct. The equations are shown below: {m/s = 2/5, (m + 8)/(s + 8) = 1/2}